OpenStudy (anonymous):

(1/z)(dz/dt) = 5 When z(1)=5 Ok so I get z=Ae^(5t) as an intermediate step, I think I got something wrong up until this point? The books answer is z=5e^(5t-5) <---- Fixed

OpenStudy (anonymous):

Wait! I fixed the problem

zepdrix (zepdrix):

Oh good you fixed it hehe.

zepdrix (zepdrix):

You are totally on the right track. Don't worry. Your next step will be to use your `initial condition` that was given, to solve for A.

OpenStudy (anonymous):

Im unsure how the -5 gets in the 5t-5 on the exponent of e

zepdrix (zepdrix):

Your A will involve an exponential of base e, they simplified the answer. You'll see after you find your A.

OpenStudy (anonymous):

ok so I set z=5, so 5=Ae^((5(1))

OpenStudy (anonymous):

I totally can't solve for A, its embarrassing but....

zepdrix (zepdrix):

hehe. Divide both sides by e^5. You'll end up with something like this,\[\large A=\frac{5}{e^5} \qquad \rightarrow \qquad A=5e^{-5}\]

OpenStudy (anonymous):

ok one second,,,

OpenStudy (anonymous):

SON of a! argh I hate hate hate this sometimes. Thanks so much without you I would have just died.

OpenStudy (anonymous):

THe book combines the e ^ to one e I get it now

zepdrix (zepdrix):

hah XD Yah it's a nice easy problem. When the book over-simplifies things, they can be difficult to recognize.